Gary Corporation produces products X, Y, and Z from a single raw material input. Budgeted data for the next month is as follows:


Product X Product Y Product Z
  Units produced 2,100       2,600       3,600     
  Per unit sales value at split-off $16       $20       $21     
  Added processing costs per unit $2       $4       $4     
  Per unit sales value if processed further $23       $23       $28     


If the cost of the raw material input is $84,000, which of the products should be processed beyond the split-off point?


Product X Product Y Product Z
  A) yes yes no
  B) yes no yes
  C) no yes no
  D) no yes yes

Option C

Option B

Option A

Option D
Answer #1
X Y Z
Sales value at split-off 16 20 21
Revenue if processed further 23-2 = 21 23-4 = 19 28-4 = 24
Incremental revenue if processed 5 (1) 3
Yes No Yes

Option B is the answer